The vapor pressure of liquid ethyl acetate, CH_3 C
Answer #1

Vapor pressure of CH3COOC2H5 is given as 40 mmHg at 282 K, that is, if the temperature falls to 282 K, the liquid will boil only if the pressure is greater than or equal to 40 mmHg. Now we have gas at 411 K and 60 mmHg. If the temperature is reduced to 282 K, its reduced pressure can be calculated according to ideal gas law at constant volume given as P1/T1 = P2/T2. Therefore 60/411 = P2/282. Therefore P2 = 41.2 mmHg.

Now since this pressure is greater than its vapor pressure at that temperature, there will be no liquid and all the CH3COOC2H5 will be in vaporized phase. No condensation will occur.